OpenStudy (anonymous):

5sqrt(-(1)/(243)) simplify

OpenStudy (watchmath):

Is it like this? \[\sqrt[5]{\frac{-1}{243}}\]

OpenStudy (anonymous):

\[5i/243\]

OpenStudy (anonymous):

The fifth root of -1/243 is the same as \[\sqrt[5]{-1}/\sqrt[5]{243}\] which comes to -1/3